# Improved multifidelity Monte Carlo estimators based on normalizing flows and dimensionality reduction techniques

### v1

@@ 3 Enhancing multifidelity estimator performance
We now describe our two methodologies for improving multifidelity Monte Carlo estimators. Our pipelines have the twofold purpose of reducing the variance of the estimator and increasing its range of applicability to problems where the original model and its low-fidelity version have a dissimilar parameterization. We recall that we consider an expensive high-fidelity model $Q^{\mathrm{HF}}\colon\mathbb{R}^{d^{\mathrm{HF}}}\to\mathbb{R}$ and a cheap low-fidelity model $Q^{\mathrm{LF}}\colon\mathbb{R}^{d^{\mathrm{LF}}}\to\mathbb{R}$, and we aim to estimate the expectation of the former model $\operatorname{\mathbb{E}}[Q^{\mathrm{HF}}(\bm{\xi}^{\mathrm{HF}})]$ under the assumption that the input parameters $\bm{\xi}^{\mathrm{HF}}\in\mathbb{R}^{d^{\mathrm{HF}}}$ and $\bm{\xi}^{\mathrm{LF}}\in\mathbb{R}^{d^{\mathrm{LF}}}$ are distributed accordingly to some probability distribution $\mu^{\mathrm{HF}}$ and $\mu^{\mathrm{LF}}$, respectively. We notice that, in the most general setting, we allow the number of input parameters $d^{\mathrm{HF}},d^{\mathrm{LF}}$ and also their distributions $\mu^{\mathrm{HF}},\mu^{\mathrm{LF}}$ to be different. Our goal is to create a shared parameterization of reduced dimension $r$, where the two models vary the most. We then assume that we want to keep the original high-fidelity model for two reasons. First, we could not have additional resources for generating new high-fidelity simulations on the shared space, and second we would like to preserve the unbiasedness of the multifidelity estimator. On the other hand, we create a modified low-fidelity model which is better correlated to the high-fidelity one. We achieve this by employing normalizing flows and dimensionality reduction techniques, in particular active subspaces for the method in Section 3.1 and autoencoders for the method in Section 3.2. We remark that the main difference between the two approaches is that autoencoders allow for nonlinear transformations which cannot be obtained from the active subspace technique. A summary of the two methodologies, which are outlined in the next two sections, is showed in Fig. 1.
@@ 3.1 Coupling MFMC with active subspaces and normalizing flows
    We notice that these approximations might not be accurate if the number $N$ of data points is not sufficiently large. On the other hand, if the surrogate models were suitably accurate, then it would make more sense to consider the neural network surrogate rather than the low fidelity model. We emphasize here that these surrogate models are employed with the only purpose of helping to find low dimensional subspaces where the models vary the most, and should not be used to approximate the models themselves for the evaluations needed to compute the multifidelity estimator.
@@ 3.2 Coupling MFMC with autoencoders and normalizing flows
    The autoencoders and the normalizing flows are trained sequentially, meaning that we first compute the best autoencoders and then train the normalizing flows on the resulting latent spaces, in order to get latent variables distributed as standard Gaussians. A similar output could also be achieved employing directly variational autoencoders. However, we decided to split the process in two steps in order to avoid a trade-off between dimensionality reduction and normalization, and obtaining the best latent space.
@@
    where both $\bm{z}_{n}^{\mathrm{HF}}$ and $\bm{z}_{n}^{\mathrm{LF}}$ are distributed accordingly to a standard Gaussian $\mathcal{N}(\bm{0},I_{r})$. We recall here that $\bm{\xi}$ stands for the original input, $\bm{x}$ is the latent variable of the autoencoder, and $\bm{z}$ is the corresponding normally distributed variable in the shared space. The next steps are analogous to the previous methodology introduced in Section 3.1. In order to increase the correlation between the high-fidelity and low-fidelity models, and without modifying the former, we construct a new low-fidelity model in which we map the input parameters of the high-fidelity model into the input parameters of the original low-fidelity model by means of the shared space. In particular, we define
@@
    Compute the normalizing flows $\mathcal{S}^{\mathrm{HF}}(\cdot;\sigma^{\mathrm{HF}})$ and $\mathcal{S}^{\mathrm{LF}}(\cdot;\sigma^{\mathrm{LF}})$ which satisfy (3.12). 3: Define the new low-fidelity model $\mathcal{Q}_{\mathrm{AE}}^{\mathrm{LF}}$ in (3.15). 4: Compute the optimal allocation $N^{\mathrm{HF}},N^{\mathrm{LF}}$ from a pilot sample. 5: Compute the estimator $\widehat{Q}_{N^{\mathrm{HF}},N^{\mathrm{LF}}}^{\mathrm{MFMC},\mathrm{AE}}$ in (3.16).
@@
    In order to train the autoencoder, it is necessary to evaluate the models $Q^{\mathrm{HF}}$ and $Q^{\mathrm{LF}}$ multiple times, and this not only is impossible for the expensive high-fidelity model, but can also be impractical for the cheaper low-fidelity model. Therefore, we propose to train surrogate models $Q_{\mathrm{NN}}^{\mathrm{HF}}$ and $Q_{\mathrm{NN}}^{\mathrm{LF}}$ for $Q^{\mathrm{HF}}$ and $Q^{\mathrm{LF}}$ based on fully connected neural networks with ReLU activation functions, analogously to what we did in Remark 3.1. We recall that these surrogate models might not be highly accurate, in particular for small amount of data, and therefore should not be used for model approximation, but only with the aim of finding nonlinear subspaces of reduced dimension during the training of the autoencoders.
@@ 3.3 A theoretical example
    Figure 3: Truncation error introduced reducing the dimensionality of the models through active subspaces and normalizing flow, for the theoretical example. The error is computed as $\left\lvert Q^{\mathrm{*F}}(x,y)-Q^{\mathrm{*F}}(\mathcal{T}^{-1}(W_{A}^{\mathrm{*F}}(W_{A}^{\mathrm{*F}})^{\top}\mathcal{T}(x,y)))\right\rvert$, where *$\mathrm{*F}$ stands for both $\mathrm{HF}$ and $\mathrm{LF}$.
@@
    $\rho_{\mathrm{AS}}=\frac{\operatorname{\mathbb{C}ov}(Q^{\mathrm{HF}}(X,Y),\mathcal{Q}_{\mathrm{AS}}^{\mathrm{LF}}(X,Y))}{\sqrt{\operatorname{\mathbb{V}ar}[Q^{\mathrm{HF}}(X,Y)]\operatorname{\mathbb{V}ar}[\mathcal{Q}_{\mathrm{AS}}^{\mathrm{LF}}(X,Y))]}}\simeq 0.98>\rho.$ (3.28)
@@
    $\rho_{\mathrm{AE}}=\frac{\operatorname{\mathbb{C}ov}(Q^{\mathrm{HF}}(X,Y),\mathcal{Q}_{\mathrm{AE}}^{\mathrm{LF}}(X,Y))}{\sqrt{\operatorname{\mathbb{V}ar}[Q^{\mathrm{HF}}(X,Y)]\operatorname{\mathbb{V}ar}[\mathcal{Q}_{\mathrm{AE}}^{\mathrm{LF}}(X,Y))]}}\simeq 0.99>\rho_{\mathrm{AS}}>\rho.$ (3.35)
@@ 4 Numerical experiments
    In this section we demonstrate the advantages of our approaches through a series of test cases. We first consider analytic functions which allow us to explore the properties of the methods, and then focus on a reaction-diffusion equation. Finally, we consider cardiovascular simulations as an example of a computationally expensive model with concrete applications and for which only a small amount of data can be available.
@@
    For all the neural networks appearing in the pipelines, after normalizing the input and output values in the interval $[-1,1]$, we perform a hyperparameter tuning for the number of layers, number of neurons per layer, learning rate, and exponential scheduler step. In particular, we tune the hyperparamters which appear in the autoencoder, the surrogate models, and the normalizing flows. The hyperparameters are optimized sequentially employing the Optuna optimization framework [1] monitoring the validation loss (20% of the dataset). In particular, we initially find the best parameters for the surrogate models, if necessary, and then use these parameters in the training process for the autoencoder. Then, once we have selected the best parameters for the autoencoder, we employ these values to compute the latent space from which we learn the normalizing flows and their corresponding hyperparameters. In the following numerical experiments, we constrain the number of layers in $\{1,\dots,4\}$, the number of neurons per layer in $\{1,\dots,16\}$, the learning rate in $[10^{-4},10^{-2}]$, and the exponential scheduler step in $[0.999,0.9999]$. Moreover, we train all the neural networks for 5000 epochs with the Adam optimizer [17], and we perform 100 independent repetitions of the entire procedure in order to illustrate its overall variability.

### v2

@@ 3 Enhancing multifidelity estimator performance
We now describe our two methodologies for improving multifidelity Monte Carlo estimators. Our pipelines have the twofold purpose of reducing the variance of the estimator and increasing its range of applicability to problems where the original model and its low-fidelity version have a dissimilar parameterization. We recall that we consider an expensive high-fidelity model $Q^{\mathrm{HF}}\colon\mathbb{R}^{d^{\mathrm{HF}}}\to\mathbb{R}$ and a cheap low-fidelity model $Q^{\mathrm{LF}}\colon\mathbb{R}^{d^{\mathrm{LF}}}\to\mathbb{R}$, and we aim to estimate the expectation of the former model $\operatorname{\mathbb{E}}[Q^{\mathrm{HF}}(\bm{\xi}^{\mathrm{HF}})]$ under the assumption that the input parameters $\bm{\xi}^{\mathrm{HF}}\in\mathbb{R}^{d^{\mathrm{HF}}}$ and $\bm{\xi}^{\mathrm{LF}}\in\mathbb{R}^{d^{\mathrm{LF}}}$ are distributed according to some probability distribution $\mu^{\mathrm{HF}}$ and $\mu^{\mathrm{LF}}$, respectively. We notice that, in the most general setting, we allow the number of input parameters $d^{\mathrm{HF}},d^{\mathrm{LF}}$ and also their distributions $\mu^{\mathrm{HF}},\mu^{\mathrm{LF}}$ to be different. Our goal is to create a shared parameterization of reduced dimension $r$, where the two models vary the most. We remark that we do not require the high-fidelity and low-fidelity models to naturally share the same lower dimensional manifold, but our main goal is finding the lower dimensional subspace of each model separately, and then create a link between them through a shared subspace. We then assume that we want to keep the original high-fidelity model for two reasons. First, we could not have additional resources for generating new high-fidelity simulations on the shared space, and second we would like to preserve the unbiasedness of the multifidelity estimator. On the other hand, we create a modified low-fidelity model which is better correlated to the high-fidelity one. We achieve this by employing normalizing flows and dimensionality reduction techniques, in particular active subspaces for the method in Section 3.1 and autoencoders for the method in Section 3.2. We remark that the main difference between the two approaches is that autoencoders allow for nonlinear transformations which cannot be obtained from the active subspace technique. A summary of the two methodologies, which are outlined in the next two sections, is showed in Fig. 1.
@@ 3.1 Coupling MFMC with active subspaces and normalizing flows
    We notice that these approximations might not be accurate if the number $N$ of data points is not sufficiently large. On the other hand, if the surrogate models were suitably accurate, then it would make more sense to consider the neural network surrogate rather than the low fidelity model. We emphasize here that these surrogate models are employed with the only purpose of helping to find low dimensional subspaces where the models vary the most, and should not be used to approximate the models themselves for the evaluations needed to compute the multifidelity estimator. Hence, even if the approximation provided by the neural networks surrogates can be quite poor, and this is a problem for identifying the best lower-dimensional subspaces, it can still be sufficient for capturing lower-dimensional manifolds that increase the correlation between the models.
@@ 3.2 Coupling MFMC with autoencoders and normalizing flows
    The autoencoders and the normalizing flows are trained sequentially, meaning that we first compute the best autoencoders and then train the normalizing flows on the resulting latent spaces, in order to get latent variables distributed as standard Gaussians. We would also like to point out that we could have used variational autoencoders which would certainly result in a more regular latent space, but without any guarantees on the distribution of the latent space to be a standard Gaussian, unlike normalizing flow where this is guaranteed by construction. Moreover, we remark that, in principle, the normalizing flows could be replaced by any other map between the latent spaces of the autoencoders of the high-fidelity and low-fidelity models, and we leave the determination of the most efficient method for this task to subsequent work.
@@
    where both $\bm{z}_{n}^{\mathrm{HF}}$ and $\bm{z}_{n}^{\mathrm{LF}}$ are distributed accordingly to a standard Gaussian $\mathcal{N}(\bm{0},I_{r})$. We recall here that $\bm{\xi}$ stands for the original input, $\bm{x}$ is the latent variable of the autoencoder, and $\bm{z}$ is the corresponding normally distributed variable in the shared space. We remark that the autoencoders do not provide a ranking of the variables in the latent dimension according to their variance contributions like the active subspace technique. Hence, if the latent space is multidimensional, we propose to select the best ordering of the components of the latent variable in terms of the estimated correlation. The next steps are analogous to the previous methodology introduced in Section 3.1. In order to increase the correlation between the high-fidelity and low-fidelity models, and without modifying the former, we construct a new low-fidelity model in which we map the input parameters of the high-fidelity model into the input parameters of the original low-fidelity model by means of the shared space. In particular, we define
@@
    Compute the normalizing flows $\mathcal{S}^{\mathrm{HF}}(\cdot;\sigma^{\mathrm{HF}})$ and $\mathcal{S}^{\mathrm{LF}}(\cdot;\sigma^{\mathrm{LF}})$ which satisfy (3.12). 3: Select the best ordering of the components of the latent variable. 4: Define the new low-fidelity model $\mathcal{Q}_{\mathrm{AE}}^{\mathrm{LF}}$ in (3.15). 5: Compute the optimal allocation $N^{\mathrm{HF}},N^{\mathrm{LF}}$ from a pilot sample. 6: Compute the estimator $\widehat{Q}_{N^{\mathrm{HF}},N^{\mathrm{LF}}}^{\mathrm{MFMC},\mathrm{AE}}$ in (3.16).
@@
    In order to train the autoencoder, it is necessary to evaluate the models $Q^{\mathrm{HF}}$ and $Q^{\mathrm{LF}}$ multiple times, and this not only is impossible for the expensive high-fidelity model, but can also be impractical for the cheaper low-fidelity model. Therefore, we propose to train surrogate models $Q_{\mathrm{NN}}^{\mathrm{HF}}$ and $Q_{\mathrm{NN}}^{\mathrm{LF}}$ for $Q^{\mathrm{HF}}$ and $Q^{\mathrm{LF}}$ based on fully connected neural networks with ReLU activation functions, analogously to what we did in Remark 3.1. We recall that these surrogate models might not be highly accurate, in particular for small amount of data, and therefore should not be used for model approximation, but only with the aim of finding nonlinear subspaces of reduced dimension during the training of the autoencoders. We also recall that even if these surrogate models are not sufficiently accurate to determine the best lower-dimensional manifolds, they can still capture nonlinear subspaces that increase the correlation.
@@ 3.2.1 A particular choice for the autoencoder
    3.3 Computational complexity analysis
    In this section we analyze the computational budget $\mathfrak{C}$ that we need to spend in order to perform our pipelines. It is clear that this additional computational cost is not required for the standard multifidelity Monte Carlo estimator. Hence, one can argue that this budget might be better invested in high-fidelity samples rather than in methodologies for increasing the correlation between the models. In the next proposition we study when it is worth spending part of the computational budget in building our modified estimators. The result is dependent on the ratio $\eta=\mathfrak{C}/\mathcal{B}$ between the cost for training the networks in the pipelines and the total computational budget, and on the improvement in term of correlation between the models.
    Proposition 3.4.
    Let $\rho$ be the initial correlation between the high-fidelity and low-fidelity models, and let $\rho_{\mathrm{AS}}$ and $\rho_{\mathrm{AE}}$ be new correlations after performing the pipelines based on active subspace and autoencoder, respectively. Assume that
    $\left\lvert\rho_{\mathrm{A*}}\right\rvert>\left\lvert\rho\right\rvert>\frac{4w}{(1+w)^{2}},$ (3.20)
    and
    $\frac{\mathfrak{C}}{\mathcal{B}}=\eta<1-\frac{\sqrt{1-\rho_{\mathrm{A*}}^{2}}+\sqrt{w\rho_{\mathrm{A*}}^{2}}}{\sqrt{1-\rho^{2}}+\sqrt{w\rho^{2}}},$ (3.21)
    where $\mathfrak{C}$ is the cost for training the networks, $\mathcal{B}$ is the total computational budget, $\mathrm{A*}$ stands for both $\mathrm{AS}$ and $\mathrm{AE}$, and $w=\mathcal{C}^{\mathrm{LF}}/\mathcal{C}^{\mathrm{HF}}$ is the cost ratio between the two fidelities. Then, it holds
    $\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{\mathcal{B}}^{\mathrm{MFMC},\mathrm{A*}}\right]<\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{\mathcal{B}}^{\mathrm{MFMC}}\right],$ (3.22)
    where both the estimators are computed assuming a total computational budget $\mathcal{B}$ and solving the optimal allocation problem.
    Proof.
    Using equation (2.8) we have
    $\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{\mathcal{B}}^{\mathrm{MFMC}}\right]=\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{\mathcal{B}}^{\mathrm{MC}}\right]\left(\sqrt{1-\rho^{2}}+\sqrt{w\rho^{2}}\right)=\frac{\operatorname{\mathbb{V}ar}\left[Q^{\mathrm{HF}}(\bm{\xi}^{\mathrm{HF}})\right]}{\mathcal{B}}\left(\sqrt{1-\rho^{2}}+\sqrt{w\rho^{2}}\right).$ (3.23)
    Moreover, since the computational budget $\mathfrak{C}=\eta\mathcal{B}$ is spent for building the modified estimators, we then employ equation (2.8) with budget $\mathcal{B}-\mathfrak{C}$, and obtain
    $\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{\mathcal{B}}^{\mathrm{MFMC},\mathrm{A*}}\right]$ $=\operatorname{\mathbb{V}ar}\left[\widehat{Q}_{\mathcal{B}-\mathfrak{C}}^{\mathrm{MC}}\right]\left(\sqrt{1-\rho_{\mathrm{A*}}^{2}}+\sqrt{w\rho_{\mathrm{A*}}^{2}}\right)$ (3.24)
    $=\frac{\operatorname{\mathbb{V}ar}\left[Q^{\mathrm{HF}}(\bm{\xi}^{\mathrm{HF}})\right]}{\mathcal{B}-\mathfrak{C}}\left(\sqrt{1-\rho_{\mathrm{A*}}^{2}}+\sqrt{w\rho_{\mathrm{A*}}^{2}}\right)$
    $=\frac{\operatorname{\mathbb{V}ar}\left[Q^{\mathrm{HF}}(\bm{\xi}^{\mathrm{HF}})\right]}{\mathcal{B}(1-\eta)}\left(\sqrt{1-\rho_{\mathrm{A*}}^{2}}+\sqrt{w\rho_{\mathrm{A*}}^{2}}\right).$
    Finally, combining equations (3.23) and (3.24) gives the desired result. ∎
    The previous result shows the condition under which we get a benefit in using our approaches. In particular, as long as the correlation increases significantly and the cost for building the modified estimators is sufficiently small compared to the total computational budget, it is better to apply our methodologies rather than using standard multifidelity Monte Carlo. This is true especially for all computationally expensive models that appear in concrete applications, such as the cardiovascular simulations in Section 4.3.
@@ 3.4 A theoretical example
    Figure 3: Truncation error introduced reducing the dimensionality of the models through active subspaces and normalizing flow, for the theoretical example. The error is computed as $\left\lvert Q^{\mathrm{*F}}(x,y)-Q^{\mathrm{*F}}(\mathcal{T}^{-1}(W_{A}^{\mathrm{*F}}(W_{A}^{\mathrm{*F}})^{\top}\mathcal{T}(x,y)))\right\rvert$, where $\mathrm{*F}$ stands for both $\mathrm{HF}$ and $\mathrm{LF}$.
@@
    $\rho_{\mathrm{AS}}=\frac{\operatorname{\mathbb{C}ov}(Q^{\mathrm{HF}}(X,Y),\mathcal{Q}_{\mathrm{AS}}^{\mathrm{LF}}(X,Y))}{\sqrt{\operatorname{\mathbb{V}ar}[Q^{\mathrm{HF}}(X,Y)]\operatorname{\mathbb{V}ar}[\mathcal{Q}_{\mathrm{AS}}^{\mathrm{LF}}(X,Y)]}}\simeq 0.98>\rho.$ (3.33)
@@
    $\rho_{\mathrm{AE}}=\frac{\operatorname{\mathbb{C}ov}(Q^{\mathrm{HF}}(X,Y),\mathcal{Q}_{\mathrm{AE}}^{\mathrm{LF}}(X,Y))}{\sqrt{\operatorname{\mathbb{V}ar}[Q^{\mathrm{HF}}(X,Y)]\operatorname{\mathbb{V}ar}[\mathcal{Q}_{\mathrm{AE}}^{\mathrm{LF}}(X,Y)]}}\simeq 0.99>\rho_{\mathrm{AS}}>\rho.$ (3.40)
@@
    Remark 3.5.
    Notice that equation (2.8) still holds true for the estimators in (3.7) and (3.16). Therefore, if we manage to increase the correlation between the high-fidelity and low-fidelity models, then we also improve the variance of the resulting estimators. For the active subspace technique, and in general for linear approaches, it is reasonable to assume that if we align the important directions of different models, then their correlation should be larger along those directions than in the original space. The intuition for this approach is provided in previous literature [12, 45]; specifically, in [45, Proposition 4.4], it is shown how the re-arrangement of the variables leads to an increased correlation in the linear case. Beyond the intuition or the quantitative analysis presented in [45] under simplifying assumptions, we also note this idea has been adopted successfully on non-trivial aerospace applications with high-dimensional inputs, see, e.g., [11, 13]. On the other hand, even if we do not have any theoretical guarantee that nonlinear lower-dimensional manifolds can increase the correlation, we expect them to behave similarly if mapped appropriately. In this work, we consider autoencoders as a nonlinear extension to linear dimensionality reduction, and the following numerical experiments highlight the possibility to improve the performance of multifidelity Monte Carlo estimators whenever a nonlinear manifold provides a more parsimonious representation than a linear subspace for the input-to-output map in at least one of the models. Moreover, we can always verify whether the new correlation, which can be estimated through a pilot sample as demonstrated in the paper, is larger than the original correlation. If this does not hold, then we can employ standard multifidelity Monte Carlo estimators, so that we can guarantee no reduction in performance apart from the negligible cost increase (compared to the models’ evaluations) of the dimension reduction and normalizing flows steps. We finally note that one could train the autoencoders for both the high-fidelity and low-fidelity models simultaneously, and including a term in the loss function that maximizes the resulting correlation.
@@ 4 Numerical experiments
    In this section we demonstrate the advantages of our approaches through a series of test cases. We first consider analytic functions which allow us to explore the properties of the methods, and then focus on a reaction-diffusion equation. Finally, we consider cardiovascular simulations as an example of a computationally expensive model with concrete applications and for which only a small amount of data can be available. We note that the dashed lines representing the “true” mean in the following plots is given by the average of the Monte Carlo estimator.
@@
    For all the neural networks appearing in the pipelines, after normalizing the input and output values in the interval $[-1,1]$, we perform a hyperparameter tuning for the number of layers, number of neurons per layer, learning rate, and exponential scheduler step. In particular, we tune the hyperparamters which appear in the autoencoder, the surrogate models, and the normalizing flows. The hyperparameters are optimized sequentially employing the Optuna optimization framework [1] monitoring the validation loss (20% of the dataset). In particular, we initially find the best parameters for the surrogate models, if necessary, and then use these parameters in the training process for the autoencoder. Then, once we have selected the best parameters for the autoencoder, we employ these values to compute the latent space from which we learn the normalizing flows and their corresponding hyperparameters. In the following numerical experiments, we constrain the number of layers in $\{1,\dots,4\}$, the number of neurons per layer in $\{1,\dots,16\}$, the learning rate in $[10^{-4},10^{-2}]$, and the exponential scheduler step in $[0.999,0.9999]$. Moreover, we train all the neural networks for 5000 epochs with the Adam optimizer [20], and we perform 100 independent repetitions of the entire procedure in order to illustrate its overall variability. The major computational cost for obtaining the networks is given by the hyperparameter tuning, which is however done as a preliminary step, before performing the multifidelity uncertainty propagation pipeline. We also notice that, as highlighted in Section 3.3, the cost for training all the networks in the pipelines is negligible with respect to the cost of high-fidelity and low-fidelity simulations, in particular for computationally expensive models, and therefore we do not include this cost in the comparison between our approaches and standard (multifidelity) Monte Carlo estimators.